# TrIM, Triangular Input Movement Systolic Array for Convolutional Neural Networks:

### v2

@@ V Design Space Evaluation
In this section, TrIM is compared to the WS-based SA and the RS-based SA through a design space evaluation. Memory accesses, throughput and number of registers are considered, following the equations introduced in Section IV. For each metric, different kernel sizes and ifmap’s sizes are considered. $K$ spans the range 3,5,7, extensively used in state-of-the-art CNNs[29, 30]. Square ifmaps ($H_{I}=W_{I}=I$) cover the range $I=[16,32,64,128,256]$.
@@ V-B Throughput
Fig. 8 illustrates the TPE for each dataflow. Fig. 8(a) refers to $K=3$, Fig. 8(b) refers to $K=5$, while Fig. 8(c) refers to $K=7$. Considering TrIM, the TPE is independent by $K$ and $I$, and reaches the peak throughput of 2 OPs/cycle/PE. This result spotlights the performance efficiency of PEs coping with the TrIM dataflow. At the PE level, TrIM behaves similarly to the WS-based SA, since weights are not moved. As a result, the TPE of WS and TrIM is the same. However, when small ifmaps are considered, the performance of WS is degraded because of the synchronization requirements of WS. This drawback is even more accentuated with larger $K$. For example, fixed $I=16$, TrIM outperforms WS by $5.3\%$, $11.8\%$, and $35.7\%$ when $K=3,5,7$, respectively.
TrIM also outperforms the TPE of RS. This because in RS each PE, after processing a row of $K$ inputs and weights, has to wait for almost the same number of cycles to allow psums to be vertically accumulated. Conversely, TrIM is an always-on dataflow, where the accumulations can be overlapped to the multiplications of new inputs and weights. The percentage improvement offered by TrIM is also reported in Fig. 8(a)-(c), reaching up to $81.8\%$.
@@ V-C Registers
Fig. 9 shows the total number of registers for each dataflow. Fig. 9(a) refers to $K=3$, Fig. 9(b) refers to $K=5$, while Fig. 9(c) refers to $K=7$. As explained in Section IV, the number of registers provides the local storing capacity of the different dataflows. Thus, simple registers, FIFOs, scratch pads and shift registers are all included in the count. The RS-based SA requires the highest number of registers, considering that memory blocks for inputs and weights must be accommodated into $K\times H_{O}$ PEs. For example, when $K=3$ and $I=256$, the registers of RS are 10$\sim$ 10$\times$ 10 × more than those required by TrIM. Since the majority of RS’ registers are memory blocks, the degradation in terms of energy is even worse.
When compared to WS, TrIM generally uses more registers if large ifmaps are considered. For example, when $I=64$ and $K=3$, TrIM requires 2.5×2.5$\times$2.5 × more registers. However, the WS-based SA predominantly makes use of FIFOs for data synchronization ($\sim 57\%$ in the example case), which are more energy demanding than SRBs in TrIM. On the contrary, when small ifmaps are considered, TrIM requires fewer registers than WS. This because SRBs are smaller than FIFOs. In fact, other than dealing with tiny ifmaps, part of the SRBs’ workload may be shifted at the PE level (when $I\leq 2K$), where PEs of adjacent rows are diagonally connected. The inversion point ($I_{IP}$) between the two trends (i.e., WS better than TrIM and viceversa) follows equation (19), obtained imposing (11) = (18), with $W_{I}=I_{IP}$:
@@ V-D Discussion
  • TrIM requires from $\times9$ to 15.6×15.6$\times$15.6 × fewer registers when compared to RS at different $K$. This is due to the simpler micro-architecture of TrIM’s PEs, which do not use memory blocks. Moreover, TrIM is advantageous when also compared to WS and for small ifmaps, considering that SRBs are smaller than synchronization FIFOs in WS. Furthermore, the larger $K$, the larger the ifmaps that can be managed by TrIM by using a lower number of registers than WS. This is advantageous with CNNs that make use of different $K$ where, usually, larger $K$ are exploited by larger ifmaps.
VI Conclusion
@@
This paper introduces TrIM, an innovative dataflow for SAs that is compatible with CNN computing. The dataflow has two levels of granularity: at the PE level, weights are kept stationary; high inputs utilization is guaranteed at the SA level, based on a triangular movement. To achieve this, $K\times K$ PEs are interconnected with each other in different directions, and local shift registers are put at the left edge of the array to assist the triangular movement of inputs. We built an analytical model to characterize the array before the physical implementation phase. Memory accesses, throughput and number of registers are subjected to in-depth design space exploration. When compared to WS dataflow, TrIM requires one order of magnitude less memory accesses considering that no data redundancy is exploited. Furthermore, TrIM requires less memory access than RS since no micro-architectural memory blocks assist data circulation. The simpler PEs in TrIM also result in 15.6×15.6$\times$15.6 × fewer registers than RS. Finally, thanks to the overlap between multiplications and accumulations, TrIM’s real throughput achieves the peak throughput of 2 OPs/cycle/PE, overcoming RS by 81.8%. In[31], we present a complete architecture dealing with TrIM and compatible with multi-dimensional convolutions for CNNs. The architecture is made reconfigurable to support different ifmap sizes and kernel sizes. Implemented onto a FPGA, the design is characterized in terms of resource utilization, performance, energy efficiency. A thorough analysis of the off-chip and on-chip memory access count is reported considering two CNN models. Comparisons with different dataflows and other FPGA implementations are reported, to highlight the advantages of the proposed TrIM-based systolic array.

### v3

@@ V Design Space Evaluation
In this section, TrIM is compared with the WS-based SA and the RS-based SA through a design space evaluation. Memory accesses, throughput, and number of registers are considered, following the equations introduced in Section IV. For each metric, different kernel sizes and ifmap’s sizes are considered. $K$ spans the range 3,5,7, extensively used in state-of-the-art CNNs[31, 32]. Square ifmaps ($H_{I}=W_{I}=I$) cover the range $I=[16,32,64,128,256]$.
@@ V-A Memory Accesses
TABLE II: Normalized Energy Estimation
$K=3$ $K=5$ $K=7$
ifmap size RS WS TrIM RS WS TrIM RS WS TrIM
$\times 16$ 16 13.9 6.9 1.2 13.9 14.1 1.7 13.9 19.1 2.3
$\times 64$ 64 15.7 8.4 1.1 15.7 22.0 1.2 15.7 40.2 1.5
$\times 256$ 256 17.5 8.9 1.0 17.5 24.2 1.1 17.5 46.7 1.1
Normalized to the energy required to read the generic ifmap once.
Table II reports the normalized energy estimation in the cases $\times 16$ 16, $\times 64$ 64, and $\times 256$ 256 ifmaps. The best results are in bold. We refer to the energy needed to read each ifmap once as a baseline, and the numbers are normalized to off-chip memory accesses, following[30]. The contribution of the RS dataflow is mainly affected by the scratch pads’ activity at the PE level. In this case, TrIM outperforms RS by up to 17.5×17.5$\times$17.5 ×. The Conv-to-GeMM conversion influences the higher energy requirements in the WS dataflow. For example, when $K=3$, WS requires from 5.8×5.8$\times$5.8 × to 8.9×8.9$\times$8.9 × more energy than TrIM.
@@ V-B Throughput
[…]
@@ V-C Registers
Fig. 9 shows the total number of registers for each dataflow. Fig. 9(a) refers to $K=3$, Fig. 9(b) refers to $K=5$, while Fig. 9(c) refers to $K=7$. As explained in Section IV, the number of registers provides the local storage capacity of the different dataflows. Thus, simple registers, FIFOs, scratch pads, and shift registers are all included in the count. The RS-based SA requires the highest number of registers, considering that memory blocks for inputs and weights must be accommodated into $K\times H_{O}$ PEs. For example, when $K=3$ and $I=256$, the registers of RS are 10$\sim$ 10$\times$ 10 × more than those required by TrIM. Since the majority of RS’ registers are memory blocks, the degradation in terms of energy is even worse.
[…]
@@ V-D Highlights
  • TrIM requires from $\times9$ to 15.6×15.6$\times$15.6 × fewer registers when compared to RS at different $K$. This is due to the simpler micro-architecture of TrIM’s PEs, which do not use memory blocks. Moreover, TrIM is advantageous when also compared to WS and for small ifmaps, considering that SRBs are smaller than synchronization FIFOs in WS. Furthermore, the larger $K$, the larger the ifmaps that can be managed by TrIM using a lower number of registers than WS. This is advantageous with CNNs that make use of different $K$ where, usually, larger $K$ are exploited by larger ifmaps.
VI Discussion
This Section provides a perspective on how TrIM can be integrated at the system level for exploitation in hardware architectures for AI. Furthermore, TrIM is contextualized in the current spectrum of dataflows based on systolic arrays, highlighting its unique advantages. Finally, we discuss how TrIM can be exploited in modern Generative AI, such as transformers and diffusion models.
Reconfigurability is one key aspect towards the system-level implementation of TrIM. This is because CNNs consist of diverse CLs, where ifmap and kernel sizes vary. In that perspective, a TrIM-based SA using $K\times K$ PEs can be made adaptable without relevant architectural refactoring. Specifically, the convolution between a $K_{E}\times K_{E}$ kernel, with $K_{E}>K$, and an $H_{I}\times W_{I}$ ifmap is realized by splitting the kernel into $K\times K$ tiles, and performing $\lceil K_{E}/K\rceil\times\lceil K_{E}/K\rceil$ convolutions, which are later summed together. In the case $(K_{E}\bmod K)\neq 0$, the kernel tiles may be padded. For example, when $K_{E}=5$ and $K=3$, four $\times 3$ 3 kernel tiles provide the equivalent $\times 5$ 5 convolution. System-level management can dynamically pause PEs’ activity for padding. From another direction, the TrIM-based SA can support different ifmap sizes by splitting the SRBs into groups. A control logic interfaces each group to the SA, according to the ifmap under processing. A SRB having length $W_{IL}-K-1$, where $W_{IL}$ is the width of the largest ifmap supported by the CNN, may be split into $N_{L}$ groups, with $N_{L}$ being the number of CLs having different ifmap sizes. The $K$ registers at the tail of each group are connected, through routing logic, to the PEs.
A special case study is implemented in[33] to accelerate CNNs like VGG-16[29] and AlexNet[3]. The architecture, consisting of 168 slices of $\times 3$ 3 PEs, has been deployed onto an AMD Zynq UltraScale+ FPGA at a clock frequency of 150 MHz. The results have confirmed the advantage in terms of memory access mitigation, up to 3×3$\times$3 × less than RS (Eyeriss chip[24]). The TrIM architecture exhibits a peak throughput of 453.6 GOPs/s, and completes the VGG-16 inference step within 78.6ms, resulting in 16×16$\times$16 × speed-up compared to RS. Comparisons with recent FPGA implementations show up to 11.9×11.9$\times$11.9 × higher energy efficiency.
TABLE III: Comparisons with Enhanced Dataflows
EOS[34] EWS[35] TrIM
Number of PEs $G\times P_{H_{O}}\times P_{W_{O}}$ $P_{M}\times P_{N}$ $P_{M}\times P_{N}\times K\times K^{\mathrm{*}}$
Ifmaps Reuse Inter-PE + ARFs Inter-PE + ARFs Inter-PE + SRBs
Weights Reuse PE Broadcast WRFs PE Stationary
Psums Reuse PE Stationary Inter-PE + PRFs Inter-PE
Local Buffers $G$ ARFs $P_{M}$ ARFs + $P_{M}\times P_{N}$ WRFs + $P_{N}$ PRFs $P_{M}\times(K-1)$ SRBs
OPs/cycle/PE (peak)∗∗ 222 222 222
Output Bandwidth∗∗ $G\times P_{H_{O}}\times P_{W_{O}}\times B$ $P_{N}\times B$ $P_{N}\times B$
We suppose $P_{M}\times P_{N}$ parallel arrays, each having $K\times K$ PEs. ∗∗Calculated.
To further steer the uniqueness of TrIM in the SAs’ spectrum, we compare its dataflow to the recent literature, where optimizations to baseline dataflows are often introduced. Relevant examples are the Enhanced Output Stationary (EOS)[34] and the Enhanced Weight Stationary (EWS)[35]. Table III summarizes the features of these dataflows compared to TrIM. As an example workload, we consider a CL with parameters $H_{O}$, $W_{O}$, $M$, $N$ and $K$ as introduced in Section II-A. The parameters $P_{H_{O}}$, $P_{W_{O}}$, $P_{M}$ and $P_{N}$ relate to the offered parallelism. $B$ is the output bit-width. EOS[34] adopts an array split into $G$ groups of $P_{H_{O}}\times P_{W_{O}}$ PEs and proposes Activation Register Files (ARFs) to improve ifmap reuse. These ARFs store redundant activations between consecutive sliding windows. EWS[35] uses $P_{M}\times P_{N}$ PEs with ARFs for redundant activations. In addition, EWS proposes Weight Register Files (WRFs) at the PE level and Psum Register Files (PRFs) at the array level to maximize the reuse of weights and psums. Figure 10 shows the register complexity ratio between EOS, EWS and TrIM. The ratio accounts for ARFs, WRFs, SRBs and PE registers for activations and weights. The analysis is based on 64 PEs. In the case of EOS/TrIM (EWS/TrIM), the register complexity ratio scales from 0.6$\sim$ 0.6$\times$ 0.6 × (1.1$\sim$ 1.1$\times$ 1.1 ×) to 1.6$\sim$ 1.6$\times$ 1.6 × (2.9$\sim$ 2.9$\times$ 2.9 ×) when example layers from the VGG-16 CNN[29] are considered. All the compared SAs exhibit a peak throughput of 2 OPs/cycle/PE. However, the real performance is affected by the required bandwidth. For instance, Table III reports the analytical expression of the output bandwidth. It can be seen that EOS delivers $G\times P_{H_{O}}\times P_{W_{O}}$ parallel convolutions, which may constraint the real throughput and the scalability of the array.
Figure 10: Register complexity ratio between the enhanced dataflows and TrIM. TrIM is considered as the baseline. The analysis refers to 64 PEs. Reuse and parallelism parameters for the enhanced dataflows follow the optimal values as in [34, 35]. Bars in pink are related to the ratio between EOS and TrIM. Dashed bars in orange relate to the ratio between EWS and TrIM. The analysis covers typical ofmap sizes as in the VGG-16 CNN[29] with $K=3$.
Other works combine the capabilities of the different dataflows towards hybrid architectures. ReDas[36] supports multiple dataflows for DNN benchmarking: complex PEs are used (with data supplied from four directions) with the support of multi-mode buffers at the array edges. TrIM, which is tailored for convolutions, simplifies the complexity of PE and buffer requirements, but preserves high data utilization: weights are reused at the PE level, ifmap are reused through triangular movement at the array level.
The effectiveness of TrIM to deal with convolutions can be also exploited in modern Generative AI: for instance, the Convolutional Vision Transformer (CvT)[12] adopts convolutional token embedding layers and convolutional projection layers to combine the invariance properties of convolutions with the global generalization and dynamic attention of transformers. Furthermore, Stable Diffusion models benefit from convolutional layers in text-to-image translation. For example, ControlNet[13] makes use of U-Net[37], which is an encoder-decoder architecture based on convolutions for down-sampling and up-sampling purposes.
VII Conclusion
@@
This paper introduces TrIM, an innovative dataflow for SAs that is compatible with CNN computing. The dataflow has two levels of granularity: at the PE level, weights are kept stationary; high input utilization is guaranteed at the SA level, based on a triangular movement. To achieve this, $K\times K$ PEs are interconnected with each other in different directions, and local shift registers are placed on the left edge of the array to assist the triangular movement of inputs. We built an analytical model to characterize the array before the physical implementation phase. Memory accesses, throughput, and number of registers are subjected to in-depth design space exploration. When compared to WS dataflow, TrIM requires one order of magnitude fewer memory accesses considering that no data redundancy is exploited. Furthermore, TrIM requires less memory access than RS since no micro-architectural memory blocks assist data circulation. The simpler PEs in TrIM also result in 15.6×15.6$\times$15.6 × fewer registers than RS. Finally, thanks to the overlap between multiplications and accumulations, TrIM’s real throughput achieves the peak throughput of 2 OPs/cycle/PE, overcoming RS by 81.8%.